OpenStudy (anonymous):

6.05] There were 150 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $405. How many Juniors bought tickets? 70 80 90 100

OpenStudy (amistre64):

j = (405 - 150(3.50))/(2.00-3.50)

OpenStudy (anonymous):

80

OpenStudy (amistre64):

j + s = 140 ; *-y jx + sy = 405 -jy -sy = -140y jx +sy = 405 ---------------- j(x-y) = 405 - 140(y) j = (405 - 140(y))/(x-y) ; where x and y are tick prices